# Redefining Machine Unlearning: A Conformal Prediction-Motivated Approach

### v1

@@ 4.1 Experimental setting
FT 0.1 $0.978({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.086})$ $0.903({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.003})$ $1.234({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.083})$ $1.317({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.173})$ $0.792({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.017})$ $0.685({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.101})$ $0.935$
@@
RL 0.1 $0.892({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.902({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.002})$ $2.639({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.488})$ $1.843({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.699})$ $0.338({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.437})$ $0.489({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.297})$ $0.971$
[…]
GA 0.1 $0.986({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.094})$ $0.900({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $1.104({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.047})$ $1.224({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.080})$ $0.894({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.119})$ $0.736({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.050})$ $0.899$
[…]
Teacher 0.1 $0.930({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.038})$ $0.902({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.002})$ $1.991({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.840})$ $1.959({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.815})$ $0.467({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.308})$ $0.460({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.326})$ $0.971$
[…]
SSD 0.1 $0.993({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.101})$ $0.897({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.003})$ $1.039({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.112})$ $1.134({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.010})$ $0.956({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.181})$ $0.791({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.005})$ $0.852$
[…]
NegGrad+ 0.1 $0.995({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.103})$ $0.848({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.052})$ $0.898({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.253})$ $1.093({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.051})$ $1.225({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.450})$ $1.287({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.501})$ $0.933$
[…]
Salun 0.1 $0.977({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.085})$ $0.924({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.024})$ $1.229({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.078})$ $1.281({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.137})$ $0.918({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.143})$ $0.884({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.097})$ $0.939$
[…]
Table 6: MIACR performance on CIFAR-10 with ResNet-18. We show the MIA value in $10\%$ random data forgetting scenario, while the MIA results in $50\%$ scenario are presented in Appendix B.
$10\%$ Forgetting $50\%$ Forgetting
Methods $\alpha$ MIACR $\uparrow$ $\hat{q}$ MIACR $\uparrow$ $\hat{q}$
RT 0.1 $0.147({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.589$ $0.201({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.570$
[…]
FT 0.1 $0.077({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.070})$ $0.627$ $0.103({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.098})$ $0.558$
[…]
RL 0.1 $0.178({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.031})$ $0.572$ $0.137({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.064})$ $0.547$
[…]
GA 0.1 $0.032({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.115})$ $0.502$ $0.055({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.146})$ $0.486$
[…]
Teacher 0.1 $0.038({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.109})$ $0.672$ $0.065({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.136})$ $0.582$
[…]
FF 0.1 $0.051({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.096})$ $0.486$ $0.089({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.112})$ $0.509$
[…]
SSD 0.1 $0.031({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.116})$ $0.511$ $0.051({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.150})$ $0.488$
[…]
NegGrad+ 0.1 $0.128({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.019})$ $0.481$ $0.109({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.092})$ $0.511$
[…]
Salun 0.1 $0.113({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.034})$ $0.681$ $0.115({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.086})$ $0.630$
[…]
Implementation Details.
For CIFAR-10/Tiny ImageNet, we randomly select $200$/$50$ data points per class ($2,000$/$10,000$ data points in total) as calibration data $\mathcal{D}_{c}$ and $\mathcal{D}_{c}^{\prime}$, respectively. The calibration data $\mathcal{D}_{c}$ does not participate in the model training or unlearning processes and is only used for calibrating the threshold $\hat{q}$, while $\mathcal{D}_{c}^{\prime}$ is used in the process of our unlearning framework to generate $\bar{q}$.
For the hyperparameter in our work, miscoverage rate $\alpha\in[0.05,0.1,0.15,0.2]$, margin parameter $\Delta=0.01$, unlearning loss weight $\lambda\in[0,0.2,0.5,1]$. Due to space limitations, more training details and experimental results under settings of $\alpha\in[0.05,0.15]$, $\lambda=0.2$ and class-wise forgetting scenario can be found in Appendix A-C.
4.2 Measure Machine Unlearning Method via New Metrics
CR Metric.
For the CR metric, we already show the results of CIFAR-10 with ResNet-18 in Tabel 4. Here, we further evaluate the unlearning performance in Tiny ImageNet with ViT under $10\%$ random data forgetting scenario by CR metric. The results of Tiny ImageNet with ViT and CIFAR-10 with ReNet18 in $50\%$ random data forgetting and class-wise forgetting scenarios can be found in Appendix B. We remark that it is computationally intractable for the FF method in Tiny ImageNet with ViT, so we omit its results. The results of Tiny ImageNet with ViT in $10\%$ random data forgetting scenario are summarized in Table 5. Our findings reveal the following insights:
For all machine unlearning methods, as $\alpha$ level increases, it results in reduced Coverage and smaller Set Size. This happens because a higher $\alpha$ loosens the conformal threshold $\hat{q}$, allowing fewer predictions to be included within the prediction set for each data point. On the contrary, the CR tends to increase with increasing $\alpha$. Although both Coverage and Set Size may decrease, Set Size often decreases more significantly. Consequently, the CR value generally becomes larger as $\alpha$ increases.
When $\alpha$ is set to $0.2$, most methods show a value of Set Size less than $1$ in both Table 4 and 5. The intuition behind it is that conformal prediction, as a static predictor, is intrinsically tied to the model’s base prediction performance and accuracy. When the model’s accuracy is significantly higher than the confidence level, conformal prediction can achieve the required coverage with ease. In fact, it can generate empty prediction sets for some data points while still meeting the target coverage. Thus, the choice of $\alpha$ is crucial. Overly high $\alpha$ values may skew evaluation results by failing to let CR accurately reflect model performance.
Notably, when $\alpha=0.2$, RL’s coverage on the forget data $\mathcal{D}_{f}$ decreases significantly compared to other methods. This results from the use of label corruption during RL’s unlearning process, which enlarges the distribution difference between calibration data and forget data. We can mitigate this gap if we know the details of specific unlearning methods, and a potential solution to this issue is discussed in Appendix B.
Regarding the forgetting quality of specific machine unlearning methods, the Teacher method achieves the strongest performance on the traditional UA metric under both evaluation criteria ➊ and ➋. However, its forgetting quality is significantly weaker when measured by the CR metric. This highlights that the CR captures critical scenarios overlooked by UA, specifically the potential retention of true labels within prediction sets for the forget data points. By explicitly addressing this phenomenon, CR ensures a more robust and reliable evaluation for unlearning quality.
@@ 4.2 Measure Machine Unlearning Method via New Metrics
Table 7: Performance of our unlearning framework. We show the performance on CIFAR-10 with ResNet-18 and Tiny ImageNet with ViT in $10\%$ random data forgetting scenario. $\lambda=0$ represents the baseline without our framework applied. It shows our framework significantly improves the forgetting quality, while preserving stable predictive performance.
@@
RT $0.1$ $8.6\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.0})$ $99.7\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.0})$ $91.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.0})$ $0.943({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.945({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $10.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.2})$ $98.3\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.4})$ $91.0\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.8})$ $0.914({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.029})$ $0.924({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.021})$ $14.0\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}5.4})$ $97.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.9})$ $90.4\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.4})$ $0.879({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.064})$ $0.912({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.033})$
$0.2$ $0.988({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.981({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.977({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.011})$ $0.976({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.005})$ $0.963({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.025})$ $0.966({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.015})$
FT $0.1$ $3.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}4.8})$ $98.1\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.6})$ $91.6\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.2})$ $0.998({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.055})$ $0.972({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.027})$ $6.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.8})$ $97.0\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.7})$ $90.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.0})$ $0.948({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.005})$ $0.924({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.021})$ $7.9\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.7})$ $96.9\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.8})$ $90.9\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.9})$ $0.940({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.003})$ $0.927({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.018})$
$0.2$ $1.000({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.012})$ $0.993({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.012})$ $0.989({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $0.974({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.007})$ $0.983({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.005})$ $0.975({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.006})$
RL $0.1$ $7.6\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.0})$ $97.4\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.3})$ $90.6\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.2})$ $0.936({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.007})$ $0.916({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.029})$ $9.7\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.1})$ $96.6\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}3.1})$ $89.4\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.4})$ $0.896({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.047})$ $0.887({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.058})$ $9.9\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.3})$ $96.9\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.8})$ $89.7\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.1})$ $0.902({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.041})$ $0.896({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.049})$
$0.2$ $0.976({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.012})$ $0.959({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.022})$ $0.964({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.024})$ $0.949({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.032})$ $0.959({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.029})$ $0.950({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.031})$

### v2

@@ 4.2 Measure Unlearning Methods via New Metrics
RT $86.92({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.091({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $82.79({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.117({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$
FT $92.00({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}5.08})$ $0.039({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.052})$ $92.92({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}10.13})$ $0.036({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.081})$
RL $74.21({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}12.71})$ $0.083({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.008})$ $61.15({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}21.64})$ $0.050({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.067})$
GA $98.80({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}11.88})$ $0.012({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.079})$ $98.86({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}16.07})$ $0.012({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.105})$
Teacher $87.24({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.32})$ $0.013({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.078})$ $93.24({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}10.45})$ $0.031({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.086})$
SSD $98.78({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}11.86})$ $0.011({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.080})$ $98.87({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}16.08})$ $0.012({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.105})$
NegGrad+ $90.30({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}3.38})$ $0.076({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.015})$ $93.82({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}11.03})$ $0.045({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.072})$
Salun $57.58({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}29.34})$ $0.055({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.036})$ $59.12({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}23.67})$ $0.044({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.073})$
SFRon $91.55({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}4.63})$ $0.017({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.074})$ $92.52({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}9.73})$ $0.017({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.100})$
MIACR Metric.
In Table 4, we show the MIACR results on CIFAR-10 under both $10\%$ and $50\%$ random data forgetting. Under evaluation criterion ➊, most methods show superior MIA and MIACR performance in the $10\%$ forgetting scenario compared to $50\%$ forgetting, because larger forget sets pose greater challenges for unlearning methods. This demonstrates that the general trend of membership leakage risk remains broadly consistent across MIA and MIACR. However, the unlearning methods considered superior under MIA metric, such as Teacher and SFRon, fail to maintain their exceptional performance when evaluated using the MIACR criterion. Similarly, under evaluation criterion ➋, Salun, which is considered optimal according to MIA metrics, does not demonstrate the best performance under MIACR evaluation. While MIA fails to accurately predict approximately $50\%$ of the forget data in Salun, MIACR can still identify membership within the prediction set.
Overall, the results show that, compared to MIACR, the existing MIA metric still leaves privacy concerns. Although MIA may fail to predict some forget data points as training members, these points can still appear in the conformal prediction set with high confidence. In contrast, MIACR more strictly controls potential membership leakage risk by measuring the probability that only non-member predictions (i.e., label ‘0’) appear in the prediction set.
4.3 Performance of Our Unlearning Framework
In this experiment, we apply RT, FT, and RL methods to our framework CPU, i.e., CPU-RT, CPU-FT, CPU-RL. Table 5 presents the results for CIFAR-10 with ResNet-18 and Tiny ImageNet with ViT in $10\%$ random data forgetting. We vary $\lambda$ in the range $[0,0.2,0.5,1]$, where $\lambda=0$ represents the baseline without our framework applied. See Table 14 in Appendix C for the results of $\lambda=1$.
@@ 4.3 Performance of Our Unlearning Framework
Table 5: Performance of our unlearning framework CPU. We show the performance on CIFAR-10 with ResNet-18 and Tiny ImageNet with ViT in $10\%$ random data forgetting. $\lambda=0$ represents the baseline without our framework applied. It shows our framework significantly improves the forgetting quality, not only across our metric but also existing metric UA, while preserving stable predictive performance.
@@
CPU-RT $8.6\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.0})$ $99.7\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.0})$ $91.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.0})$ $0.864({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.879({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $10.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.2})$ $98.3\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.4})$ $91.0\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.8})$ $0.788({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.076})$ $0.824({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.055})$ $14.0\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}5.4})$ $97.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.9})$ $90.4\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.4})$ $0.763({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.101})$ $0.825({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.054})$
CPU-FT $3.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}4.8})$ $98.1\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.6})$ $91.6\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.2})$ $0.986({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.122})$ $0.927({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.048})$ $6.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.8})$ $97.0\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.7})$ $90.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.0})$ $0.844({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.020})$ $0.829({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.050})$ $7.9\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.7})$ $96.9\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.8})$ $90.9\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.9})$ $0.853({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.011})$ $0.843({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.036})$
CPU-RL $7.6\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.0})$ $97.4\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.3})$ $90.6\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.2})$ $0.788({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.076})$ $0.796({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.083})$ $9.7\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.1})$ $96.6\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}3.1})$ $89.4\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.4})$ $0.709({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.155})$ $0.736({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.143})$ $9.9\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.3})$ $96.9\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.8})$ $89.7\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.1})$ $0.708({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.156})$ $0.731({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.148})$